### v1

Abstract
Artificial intelligence (AI) has become a crucial instrument for streamlining processes in various industries, including electrical power systems, as a result of recent digitalization. Algorithms for artificial intelligence are data-driven models that are based on statistical learning theory and are used as a tool to take use of the data that the power system and its users generate. Initially, we perform a thorough literature analysis of artificial intelligence (AI) applications related to renewable energy (RE). Next, we present a thorough analysis of renewable energy factories and assess their suitability, along with a list of the most widely used and appropriate AI algorithms. Nine AI-based strategies are identified here to assist Renewable Energy (RE) in contemporary power systems. This survey paper comprises an extensive review of the several AI techniques used for renewable energy as well as a methodical analysis of the literature for the study of various intelligent system application domains across different disciplines of renewable energy. This literature review identifies the performance and outcomes of nine different research methods by assessing them, and it aims to distill valuable insights into their strengths and limitations. This study also addressed three main topics: using AI technology for renewable power generation, utilizing AI for renewable energy forecasting, and optimizing energy systems. Additionally, it explored AI’s superiority over conventional models in controllability, data handling, cyberattack prevention, smart grid implementation, robotics- AI’s significance in shaping the future of the energy industry. Furthermore, this article outlines future directions in the integration of AI for renewable energy.
keywords:
Artificial Intelligence (AI), renewable energy, Machine Learning, Deep Learning, Power Consumption, Power System
1 Introduction
The fundamental component of financially sound, ecologically responsible, and sustainable electricity generation is renewable energy, or RE. The role that renewable energy plays in reducing climate change [1], [2], [3] is among the most compelling arguments for adopting it. When generating energy, renewable energy sources like solar, wind, and hydroelectric power emit little to no greenhouse gases, in contrast to fossil fuels [4]. When it comes to security and sustainability, renewable energy sources are superior to limited fossil fuels [5]. Natural replenishment of renewable resources, such as sunshine, wind, and water, makes them a steady and dependable energy source free from the price volatility and geopolitical unpredictability of fossil fuels [6]. By supplying power to isolated and neglected areas, renewable energy technologies provide a viable way to combat energy poverty [7]. Off-grid solar systems may provide rural communities with dependable and reasonably priced power, raising living standards and promoting economic growth. Not only is renewable energy crucial, but it is also necessary for a thriving and sustainable future. We can solve a number of issues, including as public health, economic growth, energy security, and climate change, while building a more resilient and sustainable environment for coming generations by embracing renewable energy sources and moving toward a clean energy economy [8], [9].
According to an official report by the International Energy Agency (IEA), the need for fossil fuels to generate power has begun to decline since 2019 [10], coinciding with the rise in the use of renewable energy sources to meet the world’s energy needs. In an effort to improve the efficiency of energy conversion, particularly in terms of RE generation, research on RE technologies is constantly expanding. The current global demand, particularly in developed and developing nations, necessitates the replacement of traditional electrical generation resources like fossil fuels with more sustainable energy technology [11]. Global warming and change in the climate are two harmful environmental problems brought on by fossil fuel-based energy sources. Over the past few decades, there has been an exponential increase in the greenhouse gas emissions from power generation into the environment. In order to tackle the current environmental issue, renewable energy (RE) technologies including solar, wind, hydro, biomass, geothermal, and hydrogen energies have been adopted to create electricity [12]. It is anticipated that in the future, renewable energy sources will provide all of the electricity needed to power residential areas. Data regarding Japan’s 2018 renewable energy percentage is available from the Institute for Sustainable Energy Policies (ISEP) is shown in Figure 1.
International Renewable Energy Agency (IRENA) and Renewable Energy Policy Network for the 21st Century (REN21) [13] provides the important information about Renewable Energy which is recorded in Table 1.
Table 1: Global Renewable Energy Capacity by Source (2020-2023)
Year Hydro (GW) Wind (GW) Solar (GW) Bioenergy (GW) Geothermal (GW)
2020 1305 732 707 121 14
2021 1300 780 900 130 13
2022 1350 850 1000 140 15
2023 1400 900 1100 150 16
Figure 1: The data in the pie chart shows the amount of energy that Japan’s power transmission and distribution networks transported in 2018. [14]
The global pandemic of the corona virus made it difficult to obtain current data on the world’s electrical generation in 2021 from a number of sources, as [10]’s search revealed. Figure 2 illustrates how the difference in energy generation from fossil fuels and renewable energy sources has shrunk between 2010 and 2020. From 121,531 TWh in 2010 to 136,131 TWh in 2019, the amount of fossil fuels used to generate electricity climbed somewhat, but it then declined in 2020. Conversely, the amount of energy generated by renewable sources has grown dramatically, rising from 4098 TWh in 2010 to 7,140 TWh in 2019. In contrast to the production of fossil fuels, renewable energy (RE) accounted for only 3.26–5.60
Figure 2: A comparison between the growth of renewable energy sources and fossil fuel generation (2010–2020) [10]
Figure 3 is showing forcasting that in 2025 the production of Renewable Energy will be reached to 521.95 TWh in USA. This value is incredible with compared to China, Japan, Brazil and Russia’s RE production [15].
Figure 3: TWh of data on the generation of renewable energy in the United States from 2016 to 2019 and from 2020 to 2025. Authors’ calculations serve as the source. [16]
1.1 Relevance of Renewable Energy
Renewable energy is extremely important in many aspects that are vital to sustainable development [17]. First and foremost, it plays a critical role in reducing greenhouse gas emissions, since renewable energy sources such as hydroelectric, solar, and wind power emit less greenhouse gases and provide a sustainable substitute for fossil fuels [18].In terms of the economy, the renewable energy sector supports local economies and promotes the shift to a green economy by generating employment, advancing technological innovation, and drawing investment [17]. Furthermore, the utilization of renewable energy improves public health by mitigating air and water pollution that arises from burning fossil fuels, consequently relieving respiratory ailments and other health-related issues. Most importantly, it powers socioeconomic growth by enabling fair access to electricity, especially in isolated and underprivileged areas. In addition, the shift to renewable energy drives technological improvement by promoting energy storage, grid integration, and efficiency research and development [19].
Figure 4: Sustainable eco-friendly solutions for power [20]
When discussing the issues posed by climate change, many of us immediately consider the impending energy crisis that we will almost certainly confront. Fossil fuel resources are getting harder to come by, but global CO2 emissions are still rising. Future energy requirements and security need the broad use of new renewable energy sources (shown in Figure 4). That still calls for a radical transformation in the way that scientists, students, and world leaders see renewable energy and climate change. Since renewable energy is the way of the future, we must refocus on sustainability via innovation by learning about energy, chemistry, and technology.
1.2 Renewable Energy Sector without the influence of Artificial Intelligence
Before the integration of artificial intelligence (AI) into the energy sector, renewable energy technologies such as solar, wind, hydroelectric, geothermal, and biomass were already established and evolving [21]. Developments in these technologies primarily focused on improving efficiency, reducing costs, and addressing environmental concerns. However, despite the advancements brought by AI, there were some limitations and challenges. One of the main lacking aspects was the dependency on traditional methods for data analysis and decision-making, which often lacked real-time insights and predictive capabilities [21]. Additionally, the integration of AI into renewable energy systems required significant investment in infrastructure, data collection, and training algorithms, which posed barriers to adoption, particularly in less developed regions. Moreover, there were concerns regarding data privacy, security, and the potential for AI algorithms to exacerbate biases or inequalities in energy access and distribution. Despite these challenges, the integration of AI has the potential to revolutionize the renewable energy sector by optimizing operations, improving efficiency, and accelerating the transition to a sustainable energy future. At that time researcher used to focus on improvement of renewable energy sources by means of developing tools, energy management system, finding proper solution for those which are harmful for environment. [22] this research article focuses on assessing the present status of electricity access, identifying issues and challenges, and evaluating the effectiveness of renewable energy and energy efficiency in improving electricity access in urban poor areas, specifically in the Korail slum of Dhaka (2017). This research outcome suggests that rooftop solar PV, electricity generation from solid waste, and replacing incandescent lamps with CFLs are promising approaches to enhance electricity access in the Korail slum, with significant potential for surplus electricity generation and cost savings. A separate study [23] examines the impact of environmental regulations on sustainability, taking into account factors such as research and development in renewable energy (RERD), advancements in technology (TI), and the overall economic performance (GDP). The researchers employed wavelet methods, a specialized mathematical tool, to examine the interconnections between these parameters over various time periods and in diverse manners. They discovered that increasing funding for RERD, implementing more stringent environmental regulations, and promoting the adoption of new eco-friendly technology are crucial for safeguarding the environment. Additionally, they discovered significant correlations between CO2 emissions and factors such as technological advancements and economic expansion.
1.3 Artificial Intelligence in the Renewable Energy Sector
Innovations in digital technology have the power to fundamentally alter our energy supply, commerce, and consumption. The artificial intelligence (AI) technology powers the new digitalization paradigm [24]. Intelligent software that maximizes operations and decision-making will autonomously manage the integration of energy supply, demand, and renewable sources into the power grid. AI will be essential to reaching this objective. A lot of data may be gathered by connecting the electricity grid to various devices and sensors. When combined with AI, this data can provide grid operators with fresh perspectives for more effective operation control. It gives energy providers the freedom to deftly modify supply in response to demand. AI can support distributed energy management and microgrid integration [25]. It is challenging to maintain a balance in the energy flow inside the grid when community-scale renewable energy producing units are integrated with the main grid.The quality and congestion problems may be greatly helped by the AI-powered control system. Intelligent Energy Storage (IES) and artificial intelligence combined have the potential to offer the renewable energy sector a dependable and long-lasting solution [25].
The effects of AI on the business and energy sectors are depicted in Figure 5. The influence of AI on energy enterprises across several industries is expected to exceed current expectations [24]. The influence of AI technology on various business kinds over the next five years is covered by the portion of the red line in Figure 5. The vertical axis illustrates the ”Effect of processes,” while the horizontal axis explains the ”Effect of offerings.” The ”Effect of Offerings” offers more chances to experience the influence of AI in many industries (accept or reject as desired), while the ”Effect of Processes” refers to a sequence of events or procedures followed in order to accomplish a certain objective. The majority of firms anticipate greater effects on supply chain management, production and operations, customer-focused activities, and energy information technology (IT).
Figure 5: Effects of AI in the business along with energy sectors [24]
Control solutions that make use of data-driven models of the wind-photovoltaic system employ AI and ML algorithms (Shown in Figure 6). A wind turbine’s power production may be predicted by an AI-based control algorithm based on meteorological factors like wind direction and speed [26]. Similar to this, a control system based on machine learning may be created to forecast a photovoltaic panel’s power production depending on variables like temperature and sun irradiation. The wind-PV system’s performance may then be optimized using these forecasts.
Figure 6: AI and ML-based PV-wind system control [26]
Data-driven models such as support vector machines (SVMs), k-nearest neighbors (k-NNs), regression trees, multiple linear regression (LR), gradient boosting (GB), and random forest (RF) are commonly used in the field of renewable energy (RE) prediction, leveraging machine learning (ML) techniques. Deep learning models such as long short-term memory (LSTM), gated recurrent units (GRUs), and deep neural networks (DNNs) have been used to predict power consumption and renewable energy (RE) generation across different time periods, yielding outstanding results [27], [28]. In addition, the combination of LSTM with autoencoders (AE) has produced positive outcomes [29].
1.4 Survey Framework
This research focuses on the application of AI methods in the energy industry. In order to enable researchers and readers to evaluate each other’s AI endeavors, goals, fresh, cutting-edge applications, difficulties, and future paths, this study attempts to provide a realistic baseline. Three main topics were covered: (i) utilising current developments in AI technology to create renewable power; (ii) use AI to forecast renewable energy; and (iii) optimizing energy systems. Furthermore, controllability, large data handling, cyberattack prevention, smart grid, IoT, robotics, energy efficiency optimization, predictive maintenance control, and computational efficiency were all areas in which this study examined how AI approaches surpass conventional models. The importance of AI in the next energy industry is another major topic of this study.
Figure 7: Survey framework of proposed survey research.
The survey structure of this research is organized in a strategic and logical way to cover all the important issues which is shown in the Figure 7. The rest of the paper is organized in this way.
2 AI Driven Methods For Renewable Energy
2.1 Forecasting Renewable Energy Generation using Deep Learning
Authors [30] proposes the ’AB-Net’ architecture, combining autoencoder (AE) and bidirectional long short-term memory (BiLSTM) for one-step forecasting of renewable energy (RE) generation. By acquiring data from various RE sources and deep preprocessing, the model extracts features and forecasts RE generation. Two datasets that are publicly accessible were employed in order to test and assess the efficacy of the suggested method: the wind dataset and the solar dataset. The model consists of three main steps: data acquisition from various renewable energy sources, deep preprocessing to clean and normalize the acquired data, and feature extraction and forecasting using AE and BiLSTM which is shown in Figure 8.
$h_{t}$ $=\sigma_{h}(W_{xh}X_{t}+W_{hh}h_{t-1}+b_{h})$ (1)
$y_{t}$ $=\sigma_{y}(W_{hy}h_{t}+b_{y})$ (2)
The state computation of a hidden layer in a Long Short-Term Memory (LSTM) network is explained in Equation (1). At each time step $t$, the hidden state $h_{t}$ is updated and maintained based on the previous hidden state $h_{t-1}$ and the current input $x_{t}$ to the layer. In Equation (1), $\sigma_{h}$ denotes the activation function, $W_{xh}$ is the weight matrix from the input to the hidden layer, $W_{hh}$ is the weight matrix for the recurrent connections of the hidden states, and $b_{h}$ is the bias vector for the hidden layer.
Figure 8: The design presented by the authors adheres to a well-defined framework. Data on electricity generation is acquired during Step 1. The second step entails preparing this data. During Step 3, the features are retrieved and then transmitted through the BiLSTM for the purpose of decoding. In the final step, Step 4, the model that has been trained is used to generate predictions. These predictions are then assessed using simple error measures and shown visually through graphs. [30]
AB-Net demonstrated superior performance compared to other models such as BiLSTM, CNN-BiLSTM, and an encoder-decoder (ED) in terms of mean absolute error (MAE), mean squared error (MSE), and root mean square error (RMSE). The solar dataset had mean squared error (MSE), root mean squared error (RMSE), and mean absolute error (MAE) values of 0.0106, 0.1028, and 0.0743, respectively. In contrast, the wind dataset had corresponding values of 0.0004, 0.0189, and 0.0109, respectively. This hybrid network enhances energy management efficiency by facilitating enhanced integration, trading, and control of renewable energy installations.
2.2 Big Data and Machine Learning for Smart Grid
Three layers compose the big data framework, as seen in Figure 9. The purpose of the higher, simpler layer is to store, retrieve, and perform calculations on data. Data management, sharing, and integration across many applications and domains are the responsibilities of the intermediate layer; data privacy is a major concern in this layer. The data mining platform uses data fusion technologies to preprocess data at the lowest and deepest layer.
Figure 9: A big data structure for a power utility using renewable energy [31].
The research work [31] explores the integration of big data analytics (BDA) in smart grids and renewable energy power stations, focusing on predicting system stability. A framework is developed and applied to a dataset of 60,000 instances using various machine learning methods, achieving high accuracy in predicting stability. Three machine learning models -a decision tree, random forest classifier, and conventional neural network (Deep learning) were utilized. The first model is related to the classification model, where the aim was to identify if the smart grid system will be considered stable or not. A penalized linear regression has been employed by the fourth machine learning model to estimate the stability differential equation root. Notably, the penalized linear regression model attains 96% accuracy, while convolutional neural networks (CNN) achieve 87% accuracy for classification. The main contribution lies in demonstrating the effectiveness of BDA in enhancing decision-making and system stability in smart grid environments, paving the way for future research on larger datasets and broader applications.
2.3 Using Machine Learning Models For Local Wind Power Forecasting (Neural Network Architecture and Tree-Based Approaches)
Local prosumers and consumers congregate in communities where they can collaborate or engage in competition to achieve shared goals such as reducing greenhouse gas emissions and/or power prices. Authors of the research, [32], propose data analytics modules for Renewable Energy Communities (RECs) to optimize resource usage and minimize electricity bills. It develops a day-ahead wind power forecasting algorithm using Machine Learning techniques, improving accuracy by 10% and offering representative consumption profiles for community members. The overall process is shown in Figure 10, where abnormal wind power data is recognized in the target vector and a ”forward pass” is performed to compute the actual forecast. The neural network’s parameters are changed in the backward pass after the loss function has been calculated in the third and final step.
Figure 10: Method to train neural networks when anomalous data from wind power is present. [32].
The study contributes novel methods for wind power forecasting, abnormal data handling, and generating consumption profiles, facilitating the integration of renewable energy into community grids. Deployed in a pilot REC in Belgium, the modules aim to align consumption with local generation, enhancing coordination and benefiting from advantageous pricing mechanisms. On the E-Cloud pilot project data, the ENSEMBLE model has demonstrated the greatest forecasting performance. Its output is calculated as the average of the results of four different machine learning models: random forests, gradient boosting decision trees, multilayer perceptrons, and bi-directional LSTMs. The training procedure has been modified to account for the automated detection of anomalous wind power data samples, which has significantly increased forecasting accuracy. a method that uses dynamic time warping to create community members’ representative power usage profiles.
Figure 11: Time series reveals the test set’s actual and predicted wind power for a single day. [32].
In Figure 11, the actual wind power generation (in black) and the wind power forecast (obtained with the ENSEMBLE model) as a function of time are illustrated for a random day in the test set.
2.4 Forecasting Solar Radiation using Ensemble Feature Selection and Machine Learning
Precise prediction of solar radiation is vital for the reliable operation of power systems that rely heavily on photovoltaic generation. This study [33] assesses the performance of machine learning (ML) methods in predicting solar radiation. The algorithms evaluated include Support Vector Regression (SVR), Extreme Gradient Boosting (XGBT), Categorical Boosting (CatBoost), and Voting-Average (VOA). A novel approach is introduced to choose significant variables and their previous observations, hence improving the accuracy of predicting. The ensemble feature selection approach combines Pearson’s coefficient, mutual information, random forest, and relief to choose the most related endogenous and exogenous inputs, together with their historical observation values.The suggested method for estimating solar radiation consists of five important phases, as illustrated in Figure 12. Tested on real data from Salvador, Brazil, VOA outperforms other algorithms in all prediction time horizons. Contributions include comparing ML algorithms, introducing CatBoost for solar forecasting, and proposing an ensemble feature selection method. Results show improved accuracy and highlight the importance of selecting appropriate input variables and past observations for solar radiation forecasting. For the summer and winter datasets, VOA had the faster training speed (183.93 s and 14.76 s, respectively). Being more sophisticated, VOA combines SVR, XGBT, and CatBoost. Further research is recommended to optimize parameter selection and explore applications in different locations and forecasting domains.
Figure 12: Diagram illustrating the suggested methods [33].
2.5 Optimal Hybrid AI Method (Time Delay Neural Network (TDNN)) For Smart Grid Energy Management
In order to increase the effectiveness of energy management on a smart grid like Renewable Energy Community, authors of this research [34] provide an optimum hybrid AI strategy in this work. In order to predict the future values of the energy characteristics in the community, this technique uses a Time Delay Neural Network. Afterwards, a stochastic model predictive control uses these predictions to optimize community operations using an appropriate battery energy storage system management approach. The Mean Absolute Error for solar generation, total energy consumption, and common services is 1.60 kW, 2.15 kW, and 0.30 kW, respectively, based on the outcomes of forecasts made using a public dataset with a 24-hour prediction horizon. In comparison to its rivals, the model predictive management fed with such forecasts yields the highest revenue. When compared to using the exact same system of management without taking use of forecasts from a forecasting approach, the total income is enhanced by 18.72%.
2.6 Techniques for Detecting Transient Stability Using CNN-LSTM
When a disturbance arises, electric power systems are more vulnerable to instability because they function closer to their limitations. Problems with transient stability are particularly common. [35], the research introduces a novel CNN-LSTM model (shown in Figure 13) to detect transient stability in electric power systems using historical data events. Addressing complexities in power system operation, the model considers noise, delay, loss in measurement data, line outage, and variable renewable energy (VRE) integration scenarios. Achieving over 99% accuracy in identifying stable and unstable conditions, the proposed model outperforms conventional CNN and LSTM approaches in terms of computation time, averaging 190.4 seconds. By training on synthetic data and considering various system parameters, this model demonstrates promise for enhancing transient stability detection in real-world power systems, albeit with challenges in implementation and optimization. In terms of computation time and performance, the CNN-LSTM approach outperforms the CNN and convLSTM approaches.
Figure 13: CNN-LSTM design that is being proposed [35].
3 More Application of AI in Renewable Energy Domain
The energy industry’s significant development in the use of AI, IoT, and smart grid infrastructure control has led to a study of cyber-attacks and a rapid rise in the number of solutions [36]. When it comes to load balancing renewable energy, automation is particularly crucial since it ensures grid stability by matching distributed generation to load [37]. The actions of rogue or rouge actors combined with inadequate system design, implementation, or configuration pose a threat to load balancing and the smart grid as a whole. Before a cybercatastrophy occurs, it is imperative to perform a thorough investigation about the viability and urgency of these possible dangers [37].
3.1 Implementing AI to predict renewable power
Researchers and the energy business have paid close attention to big data as well as artificial intelligence (AI) in recent years. This interest stems from two key factors: (i) growing processing power and (ii) the exponential global data expansion [38]. It is suggested to use Artificial Neural Networks (ANNs) for modeling and sizing a clean energy community (CEC) that uses energy storage devices, electric car charging stations, and a PV-wind hybrid system to fulfill building district energy demand by the authors of reserach [38]. While the second ANN is used to estimate the grid energy indication factor, the first ANN is utilized to anticipate the energy performance indicators, which include the fulfilled load fraction and the utilization factor of the energy generated. ANNs are trained on a vast amount of data in all weather scenarios, with adjustable power system configurations, and with a range of electrical demands. They use advanced models of each CEC component to anticipate the year CEC energy performance directly, without doing any system dynamic simulations. The impact of each input on each output was assessed using the Garson algorithm. Compared to those employed in ANN training, optimized ANNs have a very high prediction accuracy of CECs since they are made up of a single hidden layer with twenty neurons.
3.2 Preventing Cyberattacks in Renewable Energy Field
The clever renewable energy to cyberattacks, opening a new avenue for hackers to exploit the weaknesses in renewable energy [39]. This clearly impacts smart renewable energy’s operational and economical aspects [39]. A number of contradicting Cyber-Physical Power Systems (CPPSs) have seen moderate modifications in new kinds of power systems due to advancements in communication networks, monitoring and smart control systems, as well as the widespread usage of Internet-based structures [40]. Cyber components and power portions are typically connected in these architectures. CPPSs deal with recently developed problems like as security, vulnerability, resilience, stability, and dependability. Assessing, evaluating, and offering remedies to lessen or resolve these issues heavily rely on precise modeling techniques and investigating the interplay processes related to Smart Grids’ (SGs’) cyber-security. Many studies focused on creating reliable methods for identifying and/or defending against cyberattacks. These included studies on detection mechanisms that use federated learning techniques, perfect or imperfect false data attacks, false data-driven or data-injection attacks, and designing resilient communication networks [40].
The research article [41] introduces a novel data-driven approach for detecting false data injection (FDI) attacks in energy forecasting systems, crucial for the reliability of power systems amidst increasing reliance on renewable energy and information technology. Utilizing cross-validation, least-squares, and z-score metrics, the proposed mechanism effectively identifies intrusions without relying on system models or parameters, enhancing system resilience. Through experiments with tree-based wind power forecasting models, the approach successfully detects and removes corrupted data, restoring forecasting accuracy and generalizability.
The increasing availability of energy storage devices and the growing use of intermittent renewable energy supplies are causing significant changes in power systems [42]. These developments prompt smart-grid operators to imagine a day when microgrids would facilitate peer-to-peer energy trading, which will result in the creation of transactive energy systems. Due in major part to their remarkable durability, blockchains have attracted a lot of interest from academics and business for their possible use in decentralized TES. The research [42] investigates blockchain-based Transactive Energy Systems (TES) and introduces a novel class of attacks targeting gateways between prosumers and the system. The decentralized TES’s general architecture is depicted in Figure 14.The study formulates a comprehensive model of blockchain-based TES, analyzes various threat scenarios, and proposes attack mitigation strategies. Through experiments using GridLAB-D and a private Ethereum network, the research demonstrates that even simple attacks can manipulate market equilibria, but mitigation measures like detection and gateway switching effectively counteract these threats, ensuring the integrity and resilience of the system.
Figure 14: Infrastructure for the decentralized transactive energy system [42].
3.3 Optimizing consumption of energy employing artificial intelligence
The subject matter ”Cost Savings” via AI-powered energy efficiency in smart systems is one of the advantages. Lower operating expenses and electricity bills are directly correlated with decreased energy use. A variety of methods are included in AI technology, such as predictive modeling, data analytics, and machine learning. AI-powered systems may learn from data trends, make wise judgments, and optimize energy use in real time thanks to their capabilities. When it comes to optimizing energy usage in smart buildings, AI has several advantages. It makes it possible to gather and analyze enormous volumes of data from different sensors, meters, and Internet of things devices. One of the main causes of buildings’ energy consumption is their HVAC (heating, ventilation, and air conditioning) systems. By taking into account variables like occupancy, the state of the weather, and the need for thermal comfort, AI-based algorithms can improve HVAC operations.
[43] This study proposes a reliable method to optimize energy consumption in buildings using EnergyPlus software and a robust artificial neural network (ANN) model trained with data from a research center in Iran. Genetic algorithms are used to optimize a building’s energy usage. Highly efficient features are evaluated in the building’s energy usage. By conducting sensitivity analysis, the most influential parameters on energy consumption are identified, with the number of occupants and wall U-value being the most significant. The proposed ANN model accurately predicts energy consumption, enabling optimization through a Genetic Algorithm-based approach, resulting in a substantial 35% reduction in energy usage. Overall, the study provides valuable insights into optimizing energy consumption in buildings, offering a promising methodology for similar applications.
4 Challenges of AI adoption in the smart energy industry
The smart energy industry faces a variety of bottleneck obstacles to AI adoption, including poor or nonexistent data, problems with AI network parameters, difficulties with technical infrastructure, a shortage of skilled professionals, integration difficulties, risks, and legal and compliance concerns. For building energy systems, failure detection and diagnosis provide additional difficult issues [44]. Various studies acknowledge that inadequate information and data insecurity are two of the main problems that energy systems face [45]. The performance and dependability of the energy system are impacted by subpar sensors, controllers, and controlled devices for data estimation and energy system operation. The energy market presents additional difficulties for the power grid’s intricate correlations and strong coupling, as well as for the high data dimensionality and enormous complexity of large-scale simulation grid data [46]. For grid operations, integrating renewable energy sources like solar and wind with AI is equally challenging and complex [47].
The adoption of artificial intelligence (AI) in the smart energy industry faces several challenges. These include a lack of key AI skills among decision-makers, hindering the understanding of AI’s potential benefits [24]. Additionally, there is a shortage of practical expertise in developing reliable AI applications with real-world benefits, leading to reluctance among power companies to adopt new strategies. Outdated power system infrastructure [24]further complicates modernization efforts, with utility companies struggling to manage vast amounts of dispersed and disorganized data. Economic pressures also impede the integration of innovative energy technologies due to high costs and resource requirements. Furthermore, decentralization and diversification of energy supply, coupled with emerging AI technologies, present complex challenges for energy production and distribution worldwide. Dependency on cellular technologies limits AI’s potential [24], especially in developing economies. Moreover, AI-based applications often act as black boxes for consumers, raising concerns about transparency and security, particularly in the face of increasing cyber threats in the digital energy infrastructure. Addressing these challenges is essential to unlock the full potential of AI in revolutionizing the smart energy industry.
5 Current Challenges and Future Research Directions
The integration of artificial intelligence (AI) in renewable energy research poses various challenges and opens avenues for future exploration. Firstly, the complexity and cost of AI algorithms used in optimization, design, and management of renewable energy systems need to be addressed. Simplifying and making these models cost-effective is crucial for widespread adoption and scalability.
Addressing real-time issues such as automatically controlling load demand for wind and solar power farms under extreme weather conditions, enhancing the efficiency of geothermal plants using IoT and AI, automating circuit breakers and relays under fault conditions, and preventing/detecting the impact of lightning on transmission lines remain critical challenges.
Regulators and policymakers must recognize AI’s systemic impact on reducing greenhouse gas emissions and ensure flexible policies to facilitate the rapid deployment of AI technologies in the energy sector. Incorporating ethical implications of AI in corporate social responsibility standards is essential for responsible AI deployment. Establishing participatory mechanisms to understand stakeholder expectations and limitations in AI application cases is imperative. Building AI operating systems that accurately identify energy efficiency and cost savings for consumers and businesses will be invaluable.
Furthermore, enhancing transparency and accessibility of system data, protecting AI-based models from security risks, and integrating AI with other technologies like robotics, sensors, and IoT for optimization are key areas for future research. With distributed energy resources and microgrids, artificial intelligence (AI) may be a key component of the grid’s integration of renewable energy sources and promote operational independence. AI may be introduced to developing power markets with the help of international financial institutions, enabling a sustainable energy transition.
Last but not least, realizing AI’s promise in energy systems calls for a multifaceted strategy, assessing the viability of integrating AI, finding interesting AI methodologies, and offering practical insights for improving performance while pleasing a range of multidisciplinary audiences. In order to solve the difficulties posed by global climate change and achieve sustainable energy transitions, it is imperative that comprehensive frameworks and techniques for the efficient integration of AI in energy systems be developed.
From revolutionizing medical diagnoses [48], [49], [50], [51], [52] to enhancing personalized learning in education, and optimizing transportation networks for efficiency, artificial intelligence has been instrumental across various sectors [53], [54]. Now, with recent advancements, the power sector is harnessing AI’s potential to improve grid management, forecast energy demands, and enhance overall operational efficiency. Researchers anticipate that major gains in energy efficiency and sustainability will result from the integration of artificial intelligence (AI) with renewable energy sources. Artificial intelligence (AI) has the potential to increase forecasting accuracy, stabilize the grid, optimize renewable energy systems, and provide more intelligent energy management options. Given AI’s capacity for massive data analysis and real-time decision-making, we see a future in which renewable energy sources will predominate in supplying the world’s energy needs while minimizing its negative environmental effects. A more robust and sustainable energy environment is anticipated as a result of the convergence of AI with renewable energy, opening the door to a cleaner and greener future.
Table 2: An overview of this survey article that includes the methodology and results of previous studies
Reference Year Methodology Model Data Source Research Outcome
[30] 2021 Used deep learning (DL) method AB-Net for forecasting of renewable energy generation Solar Dataset [55], Wind Dataset [56] AB-Net approach provides MSE 0.0004, RMSE 0.0189.
[31] 2022 Used decision tree, random forest, and conventional neural network, linear regression for smart grid management Smart Grid data [57] Linear regression model attains 96% accuracy.
[32] 2021 Developed a day-ahead wind power forecasting algorithm using Machine Learning algoriths (random forests, gradient boosting decision trees, multilayer perceptrons, and bi-directional LSTMs) to minimize electricity bills Meteo Data (Wind speed, pressure, temperature) [58], [32] Ensemble model provided improved RMSE 2327 kW.
[33] 2023 Extreme Gradient Boosting (XGBT), Support Vector Regression (SVR), Categorical Boosting (CatBoost), and Voting-Average (VOA), which combines SVR, XGBT, and CatBoost, are the machine learning techniques that are employed for forecasting of solar radiation Brazilian National Institute of Meteorology website (INMET) [59] The training speed of VOA was greater than that of XGBT. Winter-VOA forecasting accuracy was 0.2417 MAE, whereas summer-VOA forecasting accuracy was 0.2877 MAE.
[34] 2022 Time Delay Neural Network for energy management Data of PV generation plant [60] Total income is enhancedby 18.72%
[35] 2022 CNN-LSTM model to detect transient stability in electric power systems synthetic dataset generated from the DigSILENT PowerFactory simulation, available at: [61] Discovery of transient stability exceeded 99%.
[38] 2022 Artificial Neural Networks (ANNs) for electric vehicle (EV) charging station and wind-photovoltaic battery HRES performance assessment The computation of the eight dimensionless input data requires the input climate data from the 48 locales and the load data from the three building districts under consideration [38] The R2 values for the SLF, UF, and GEIF forecasts are, in order, 0.952, 0.925, and 0.942.
[42] 2021 explored TES based on blockchain and presented a new category of cyberattacks Historical data on the prosumers’ previous bids Threats reduced by switching the gateway and using detection.
[43] 2020 Genetic algorithms are used to optimize a building’s energy usage. Highly efficient features are evaluated in the building’s energy usage. After being gathered, weather information was loaded into the Metronome program (”National and Local Weather Radar, Daily Forecast, Hurricane and information from The Weather Channel and weather.com,” n.d.) Garson’s method demonstrates the ANN model’s excellent accuracy of 99.9% when it comes to remodeling.
6 Conclusion
An extensive review of the use of artificial intelligence (AI) approaches in renewable energy (RE) systems is given in our survey article. Nine AI-based techniques for modern power systems—which address topics including renewable power production, energy forecasting, and system optimization—are found by means of a comprehensive literature study. The reviewed research demonstrate how several AI methods, like as neural networks, decision trees, random forests, and deep learning, are successful in tackling a range of RE management concerns.
The surveyed research demonstrates the potential of AI in enhancing controllability, data handling, cyberattack prevention, and smart grid implementation within the energy industry. Furthermore, the superiority of AI models over conventional approaches is highlighted, particularly in terms of accuracy, efficiency, and adaptability to complex scenarios. Future directions for the integration of AI in renewable energy systems are outlined, emphasizing the need for continued research and development in this rapidly evolving field. Key areas for future exploration include the refinement of AI algorithms, the integration of AI with other emerging technologies such as blockchain and IoT, and the establishment of regulatory frameworks to support AI adoption in the energy sector.
Overall, this survey paper contributes valuable insights into the state-of-the-art AI applications in renewable energy and provides a roadmap for future research endeavors aimed at leveraging AI for sustainable energy transition and environmental conservation.
Declarations
Funding
A fraction of the financial support for this research is provided by DOEd Grant P116Z220008 (1) and NSF Grant No. 2306109. The author(s) presents all viewpoints, discoveries, and conclusions; the sponsor(s) may not necessarily endorse them.
Conflict of interest
The authors have no conflicts of interest to declare.
Ethics approval
This review paper does not include any original research that involves human beings, animal subjects, or sensitive data.
Consent to participate
The papers and data included in this review were obtained from publically accessible and previously published literature.
Consent for publication
All authors mentioned in this review study have examined and endorsed the final work for submission to the Machine Learning journal. Every author has made substantial contributions to the research, writing, and revision of the manuscript. All writers have thoroughly reviewed and approved the content included in this review, and they have given their explicit agreement for its publication.
Availability of data and materials
The data and materials utilized are available in the references of this study.
Code availability
The review is conducted by thoroughly analyzing publically accessible literature, and any mentions of particular approaches or algorithms are properly credited to their original sources. We have made diligent efforts to include explicit citations and references to facilitate readers in locating the source works for additional scrutiny.
Authors’ contributions
Conceptualization: [Parag Biswas], [Angona Biswas]; Methodology: [Abdur Rashid], [Angona Biswas]; Formal analysis and investigation: [Angona Biswas], [Abdullah Al Nasim]; Writing - original draft preparation: [Angona Biswas][Abdullah Al Nasim], [Abdur Rashid],[Parag Biswas]; Writing - review and editing: [Abdur Rashid]; Funding acquisition: [Kishor Datta Gupta]; Resources: [Abdur Rashid], [Parag Biswas]; Supervision: [Kishor Datta Gupta]